OpenStudy (anonymous):

H-E-L-P so frustrating!! Determine the extrema of below on the given interval f(x)=5x^3-61x^2+16x+3 (a) on [0,4] The minimum is ?? and the maximum is ?? (b) on [-9,9] The minimum is ?? and the maximum is ?? please solve so i can know which answer i got wrong.

OpenStudy (anonymous):

can someone please tell me the correct answer to this. it seems like one of my answers is incorrect.

OpenStudy (anonymous):

check \(f(0), f(4)\) and also find the critical point and find \(f\) of that number

OpenStudy (anonymous):

a) on [0,4] The minimum is -589 and the maximum is 4.06074 (b) on [-9,9] The minimum is -8727 and the maximum is -1149

OpenStudy (anonymous):

the derivative is \( 15 x^2-122 x+16\)

OpenStudy (anonymous):

which, by some miracle factors as \((x-8) (15 x-2)\)

OpenStudy (anonymous):

zeros of the derivative are \(\frac{2}{15}\) and \(8\)

OpenStudy (anonymous):

so \(\frac{2}{15}\) is the \(x\) coordinate of the local max and \(8\) is the \(x\) coordinate of the local min

OpenStudy (anonymous):

so for the intervals [0,4] max is: 2/15 and 8 is the min? what about [-9,9]

OpenStudy (anonymous):

i just want to make sure because it submitted the answer and it keeps saying its wrong.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

any one! this is important!

OpenStudy (anonymous):

never mind! i found it. thanks to those who actually helped.

OpenStudy (anonymous):

f'(x)=10x^2 - 122x +16 0=10x^2 - 122x + 16 (just working here because I don't have paper handy)

OpenStudy (anonymous):

careful here the max is the \(y\) value not the \(x\) value

OpenStudy (anonymous):

just finding critical points

OpenStudy (anonymous):

oh, sorry you already found them

OpenStudy (anonymous):

8 is not in the interval \([0,4]\) for that one, you need to check \(f(0), f(4), f(\frac{2}{15})\) the largest is the max and the smallest is the min

OpenStudy (anonymous):

for \([-9,9]\) you need to check \[f(-9),f(\frac{2}{15}), f(8), f(9)\]

OpenStudy (anonymous):

got it. thank you!

OpenStudy (anonymous):

yw